A boat moves through the water with two forces acting on it. One is a 1,950-N forward push by the water on the propeller, and the other is a 1,300-N resistive force due to the water around the bow.

(a) What is the acceleration of the 1,100-kg boat? m/s 2

(b) If it starts from rest, how far will the boat move in 20.0 s?  m

(c) What will its velocity be at the end of that time? m/s
Answer #1

Ans=

1. Acceleration=

Fnet = ma

1950 - 1300 = 1100a

a = 0.6 m/s2  

2. Distance travelled by boat, intially at rest = u=0

S = 0.5at2  

S = 0.5*0.6*400 = 120m

3) Velocity at the end of time

v2 = 2*9.8*120

v = 49 m/s